Can someone help me with this equation I am having some difficulty and tried multiple ways .This is a Linear equation with quadratics.
x^2+3=y
y=5-x

Answer by scott8148(6628)

substituting ___ x^2 + 3 = 5 - x ___ x^2 + x - 2 = 0

factoring ___ (x + 2)(x - 1) = 0

there are two sets of solutions ___ the line intersects the parabola at two points